OpenStudy (imtant):

How to integrate 4 sin 2t dt?

OpenStudy (psymon):

\[\int\limits_{}^{}4\sin2t dt\] let u = 2t, therefore du = 2dt dt = du/2 \[\int\limits_{}^{}4sinu*\frac{ du }{ 2 }= 2\int\limits_{}^{}sinu\] the integral of sinu is -cosu, meaning our answer is: \[-2cosu + C = -2\cos(2t) + C\] At the end I just replaced u with what I originally said it was.

OpenStudy (imtant):

Ok I got it. Thanks Psymon. :)

OpenStudy (psymon):

Okay, awesome.

OpenStudy (imtant):

How about integrate -y(1+2t^2) dt?

OpenStudy (psymon):

-y and t in the same equation?

OpenStudy (imtant):

Ops sorry.Wrong typing. It's -(1+2t^2) divide t dt.

OpenStudy (psymon):

So: \[\frac{ -(1+2t^{2}) }{ t }\] ?

OpenStudy (imtant):

Yes it is.

OpenStudy (psymon):

Gotcha. Well, split it up into two fractions: \[\frac{ -(1+2t^{2}) }{ t }= -\frac{ 1 }{ t }-\frac{ 2t^{2} }{ t } \implies -\int\limits_{}^{}\frac{ 1 }{ t }-2\int\limits_{}^{}t\]As long as you know how to integrate those now youre good.

OpenStudy (imtant):

Ok so I will get - In t - 2t^2 /2.Right?